OpenStudy (anonymous):

is sin10x=10sinx? give a counter example if its false.

OpenStudy (anonymous):

pick any number for x and you will see. well don't pick x = 0

myininaya (myininaya):

no if x=pi/2 then sin(pi/2)=1 10sin(pi/2)=10(1)=10 sin(10*pi/2)=sin(5pi)=0

OpenStudy (anonymous):

maybe pick \[x=\frac{\pi}{2}\] then get \[\sin(10\times \frac{\pi}{2})=\sin(5\pi)=0\]

OpenStudy (anonymous):

whereas \[10\sin(\frac{\pi}{2})=10\times 1=10\]

OpenStudy (anonymous):

oh look, we used the same exact example!

OpenStudy (anonymous):

ok i admit i copied.

myininaya (myininaya):

lol

myininaya (myininaya):

well pi/2 was easier choice to use

myininaya (myininaya):

easiest*

OpenStudy (anonymous):

next time i will pick \[\frac{\pi}{12}\]

myininaya (myininaya):

lol

OpenStudy (anonymous):

you teach tuesday?

myininaya (myininaya):

yes

OpenStudy (anonymous):

have fun!

myininaya (myininaya):

it will be a blast im sure

OpenStudy (anonymous):

thank you so much(:

OpenStudy (anonymous):

Can you detect a possible danger in determining the domain of a function solely by analyzing the graph generated by a graphing utility?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

especially if your graph has a "hole" in it. here is an example http://www.wolframalpha.com/input/?i=y%3D%28x^2-4%29%2F%28x-2%29

OpenStudy (anonymous):

the function is not defined at x = 2 but the graphing utility will not pick it up

OpenStudy (anonymous):

I need help getting started on writing a short paragraph about the importance of examining a function analytically as well as graphically.